OpenStudy (anonymous):

Pleaseeeee helppp meeee :( Which is correct about the inflection point for the function f(x)= e^x - 2x^2 inflection point is at x=0 '' '' '' is at x=1.39 '' '' '' is at x=1.5 '' '' '' is at x=1.69

hero (hero):

You don't have a graphing calculator?

OpenStudy (anonymous):

no:(

OpenStudy (anonymous):

you need to take the second derivative first, that would be: \[f(x) = e^x-2x^2 \Rightarrow f'(x) = e^x-4x \Rightarrow f''(x)=e^x-4\] Setting the second derivative equal to zero we obtain: \[0 = e^x-4 \Rightarrow e^x = 4 \Rightarrow x = \ln(4)\] According to wolfram alpha, ln(4) = 1.386

OpenStudy (anonymous):

You are amazing!!!